OpenStudy (anonymous):

At the start of the basketball game, the referee tosses a ball for a jump ball. The equation that models the height of the ball is h(t) = -16t2 + 24t + 5 During what time interval is the height of the ball above 9 feet? (Round your answer to the nearest tenth of a second.)

OpenStudy (anonymous):

did this exact problem before yes?

OpenStudy (anonymous):

-16t^2+24t+5=9 t = 1/4 (3+sqrt(5)) t=1.31 seconds

OpenStudy (anonymous):

This is a problem that tests if you can use the quadratic formula. Start by setting h(t)=9. This gives you -16t^2+24t-4=0.This gives you the solutions t = 0.191 and t = 1.31. You know the graph of this equation is a parabola facing down, and thus the interval must be in between these 2 solutions. So your interval is (0.2, 1.3).